OpenStudy (anonymous):

find the real solution of the equation. \[\sqrt[5]{1-x}=-2\] the answer is 33. show me how and why please.

OpenStudy (anonymous):

raise both sides to 5th power

OpenStudy (anonymous):

\[\sqrt[5]{1-x}=-2\]

OpenStudy (anonymous):

you should get: 1-x=-32, then solve for x

OpenStudy (anonymous):

x = 33